OpenStudy (hkmiu):

what basic trigonometric identity would you use to verify that sinx cosx tanx=1-cos^2x

OpenStudy (tkhunny):

First one that came to my mind: \(\tan(x) = \dfrac{\sin(x)}{\cos(x)}\).